OpenStudy (anonymous):

Use implicit diff. to find dy/dx for x^2 + y^2 = 4

OpenStudy (anonymous):

2x + ( 2y *dy/dx) =1 (2y * dy/dx) = 1- 2x dy/dx = (1-2x )/2